OpenStudy (thefluffmuncher):

Look at the cylinder shown. The cylinder has a radius of 3 centimeters and a height of 8 centimeters. What is the volume of the cylinder? A 24π cubic centimeters B 66π cubic centimeters C 54π cubic centimeters D 72π cubic centimeter

OpenStudy (thefluffmuncher):

@Bella5

OpenStudy (anonymous):

v is pi r square h therefore v = 22/7 x 3 square x 8 Hope it helped

OpenStudy (wolf1728):

Volume = PI * r² * height Volume = 226.19 and that is not one of the choices

OpenStudy (thefluffmuncher):

yea i know

OpenStudy (wolf1728):

The answer is D 72 PI

OpenStudy (thefluffmuncher):

ty

OpenStudy (thefluffmuncher):

i got a couple more if u got the time

OpenStudy (wolf1728):

u r welcome (I hate those stupid ways of expressing an answer. DON'T say 226.19 - say 72 PI ARRGGHH)

OpenStudy (wolf1728):

I've got the time

OpenStudy (thefluffmuncher):

this one i was confused by

OpenStudy (thefluffmuncher):

Evaluate: 6!

OpenStudy (thefluffmuncher):

A 21 B 720 C 30 D 36

OpenStudy (wolf1728):

That is 6 factorial and it equals 6*5*4*3*2*1 = 720 answer is B

OpenStudy (thefluffmuncher):

oh ty so much now i know how to do them

OpenStudy (wolf1728):

didn't they explain factorials to you?

OpenStudy (thefluffmuncher):

no

OpenStudy (wolf1728):

Just a quick lesson. Take the word "strong" 6 letters no repeated letters. How many ways can you arrange the 6 letters? 6! or 720

OpenStudy (thefluffmuncher):

Sonal is making a rectangular banner for her school. She first makes a scale drawing of the banner on a piece of paper before making the larger banner. Her scale drawing is shown. What will be the actual dimensions of the banner? A 12 feet by 30 feet B 9 feet by 22.5 feet C 6 inches by 15 inches D 3 inches by 7.5 inches

OpenStudy (thefluffmuncher):

you just thought me how to do factorials and im grateful for this!!

OpenStudy (wolf1728):

thanks- by the way - in the latest problem you are going to post a graphic right?

OpenStudy (thefluffmuncher):

im thinking its b

OpenStudy (wolf1728):

1 inch = 1.5 feet so the banner is 9 x 22.5 feet You're right - the answer is B

OpenStudy (thefluffmuncher):

Look at this shape. What is the area of the shape? https://acuityk12.ctb.com/images/atom11400_11499/11401/probs/11401_15392_91682_P_MATH_2.gif A 18π + 84 square units B 18π + 66 square units C 72π + 66 square units D 36π + 66 square units

OpenStudy (wolf1728):

Area of Circle = PI*r²/2 = PI 18 Area of rectangle + triangle = 48+18 = 66 Total = PI 18 + 66 or 18 PI + 66 it is B

OpenStudy (thefluffmuncher):

Sami is selling lemonade. He has 9 1/3 gallons of water and 2 1/6 gallons of lemonade mix. He wants to use the same amount of lemonade mix for each gallon of water. Which of these is the amount of mix Sami should use for each gallon of water?

OpenStudy (thefluffmuncher):

A.1 5/13 B.20 2/9 C.13/26 D.13/18

OpenStudy (wolf1728):

It seems 2 + 1/6 gallons of mix are to be mixed with 9 + 1/3 gallons of water. 2.16666666666666 / 9.3333333333 = 0.2321428571 gallons of mix for every gallon of water and that answer isn't there is it?

OpenStudy (thefluffmuncher):

yea i dont see it

OpenStudy (thefluffmuncher):

@rvc

OpenStudy (wolf1728):

The trouble with fractions typed on a computer is that they can be mistaken for something else. Is 1 5/13 1 + 5/13 or 15/13?

OpenStudy (wolf1728):

By the way the answer I got 0.2321428571 equals 13/56

OpenStudy (thefluffmuncher):

1 5/13 the one is whole number

OpenStudy (wolf1728):

Is C 13/26? - if it is 13/56 then that is the answer.

OpenStudy (thefluffmuncher):

yea ill just go with that

OpenStudy (thefluffmuncher):

How many different arrangements are possible for 5 people standing side-by-side in a line?

OpenStudy (thefluffmuncher):

120?

OpenStudy (wolf1728):

I'd say that would be it - 120 (or 5!)

OpenStudy (thefluffmuncher):

its c its 13/56 not 13/26

OpenStudy (thefluffmuncher):

that was a derp on my part

OpenStudy (wolf1728):

So C was a typo - it sgould be 13/56?? All right !!!

OpenStudy (thefluffmuncher):

Which of these best describes a triangle with angle measures of 130°, 20°, and 30°?

OpenStudy (thefluffmuncher):

A right triangle B acute triangle C obtuse triangle D equilateral triangle

OpenStudy (wolf1728):

C - obtuse triangle it has one angle greater than 90° Here's a good web page about triangles written by ... me. http://1728.org/triang.htm

OpenStudy (thefluffmuncher):

Line l is parallel to line m, as shown. What is the measure of angle w?

OpenStudy (wolf1728):

Angle W = 145°

OpenStudy (thefluffmuncher):

Hannah earned 78 dollars in 6 hours. What is the ratio of money to time?

OpenStudy (thefluffmuncher):

A 72 dollars/hour B 13 dollars/hour C 72 hours/dollar D 13 hours/dollar

OpenStudy (thefluffmuncher):

thinking b

OpenStudy (wolf1728):

78 dollars/6 hours = 13 dollars per hour so it is B

OpenStudy (thefluffmuncher):

Look at this compound inequality. Which graph represents the solution to the compound inequality?

OpenStudy (wolf1728):

It would be graph D

OpenStudy (thefluffmuncher):

Juan is planting a vegetable garden in his backyard. In the garden, 5 out of the total 25 plants are tomato plants. What percentage of the plants are tomato plants?

OpenStudy (thefluffmuncher):

A 80% B 30% C 500% D 20%

OpenStudy (thefluffmuncher):

its d

OpenStudy (wolf1728):

% = (5/25) x 100 = .2 * 100 = 20% yes it is D

OpenStudy (thefluffmuncher):

A working replica of a steam locomotive can travel 2 kilometers using the energy from burning 9 kilograms of coal. In 12 hours the locomotive travels 18 kilometers. How many hours will the locomotive be able to travel using the energy from burning 27 kilograms of coal?

OpenStudy (thefluffmuncher):

A 18 B 4 C 162 D 46

OpenStudy (wolf1728):

Well, it's 3:00 am here - I'll have to get going - sorry. Anyway good luck.

OpenStudy (thefluffmuncher):

ok ty for everything u were tremenduos help

OpenStudy (wolf1728):

thanks - ok see ya